A 20-kg child on a swing can make a maximum angle of 400 with respect to the vertical. , what is the child’s maximum speed if the chains of the swing have a length of 4 m?


1
Expert's answer
2021-11-21T17:26:47-0500

"h=4-4\cdot\cos40°=0.94\ (m)"


"mgh=mv^2/2\to v=\sqrt{2gh}=\sqrt{2\cdot9.8\cdot0.94}=4.3\ (m/s)"